# Targeting Misalignment: A Conflict-Aware Framework for Reward-Model-based LLM Alignment

### v1

@@ Introduction
Large Language Models (LLMs) have been undergoing rapid evolution (guo2025deepseek; pichai2024our; achiam2023gpt), and have demonstrated remarkable capabilities across a wide range of tasks, including following instructions (chung2024scaling; ouyang2022training), performing complex reasoning (anil2023palm; wei2022chain), and code generation (gao2023pal; ni2024l2ceval). A key method for aligning these models with human intent is to fine-tune them using reward signals. A standard paradigm for this is Reinforcement Learning from Human Feedback (RLHF), where a pre-trained base policy $\pi_{\text{base}}$ is optimized using a reward model $r$ that approximates human preferences across dimensions such as helpfulness, safety, or formatting (christiano2017deep; chen2024mj; ziegler2019fine). However, this approach rests on the assumption that the reward model accurately captures the true supervision signal. In practice, this assumption is frequently violated due to annotation noise, limited or biased human data (jeon2020reward; sadigh2017active). As a result, the learned reward function becomes an imperfect proxy ($r_{\text{proxy}}$) of the true objective, which can lead to unintended behavior of the policy or sometimes serious consequences, such as reward over-optimization (coste2023reward; moskovitz2023confronting) or reward hacking (laidlaw2024correlated; wang2025beyond; denison2024sycophancy; everitt2021reward; pan2022effects), where models optimize the proxy reward while diverging significantly from the actual intent.
In this work, we investigate how such misalignments manifest in practice and how we might systematically surface and mitigate their effects. Specifically, we consider a realistic scenario where we start from a strong base model $\pi_{\text{base}}$, already proficient on some tasks, and fine-tune it using a proxy reward $r_{\text{proxy}}$ trained on noisy or limited preference data from a related but distinct task. This setup is particularly relevant in practical settings where powerful models like GPT-4 (achiam2023gpt) already demonstrate strong general capabilities, and fine-tuning is applied to adapt them to niche or sensitive objectives. In such cases, we want to avoid degrading the base model’s performance or introducing unintended behaviors due to flawed or imprecise proxy supervision. Prior work has attempted to mitigate these effects either by regularizing the fine-tuning process, such as through KL-penalty (liu2020learning), or by enhancing the proxy reward model itself (shen2024improving; rame2024warm). While effective, these approaches often lack insight into why such solutions work or where exactly the proxy model may be going wrong.
@@
Aligning LLMs with human intent often follows the reward-model-based alignment paradigm, also known for RLHF (christiano2017deep; stiennon2020learning), which typically consists of three major stages: (1) supervised fine-tuning, (2) reward modeling, and (3) reinforcement learning.
@@
Reward Modeling: To emulate human preferences, a reward model $r(x,y)$ is trained to score responses given a prompt $x$. A standard approach is to learn from pairwise human comparisons using the Bradley-Terry (BT) model (bradley1952rank), which defines the probability of preferring response $y_{1}$ over $y_{2}$ as: $p^{*}(y_{1}\succ y_{2}|x)=\frac{\exp(r(x,y_{1}))}{\exp(r(x,y_{1}))+\exp(r(x,y_{2}))}$. Here, $y_{1}\succ y_{2}$ denotes $y_{1}$ is preferred and $y_{2}$ is dispreferred. Given a dataset $D=\{x^{i},y_{\omega}^{i}\succ y_{l}^{i}\}_{i=1}^{N}$ consisting of prompt-response pairs with human-labeled preferences, the reward function can be trained by minimizing the following logistic loss:
@@
This objective is typically optimized using proximal policy optimization (PPO) (schulman2017proximal).
@@
Proxy-Policy Alignment Conflict Score (PACS). A well-known observation in RLHF is that the optimal policy $\pi_{r}$ under a reward function $r$ for Equation 2 satisfies (peng2019advantage; peters2007reinforcement; rafailov2023direct):
@@
Kendall-Tau Distance (K-T Distance). While PACS provides a localized conflict score, we also seek to assess the overall alignment between the base model and the proxy reward across a distribution of completions. To this end, we adopt the Kendall-Tau Distance (abdi2007kendall), a standard measure of disagreement between two ranked lists.
@@
Building on the earlier insights that proxy-policy conflicts serve as valuable signals for additional supervision, and leveraging the defined conflict detection metrics, we now propose a practical algorithm SHF-CAS: Selective Human-in-the-Loop Feedback via Conflict-Aware Sampling. The complete procedure is described in Algorithm 1. It starts with identifying candidate QA pairs that exhibit high conflicts using the proposed PACS and K-T Distance metrics. These selected examples are then prioritized for human feedback. In particular, to make this approach resource-efficient, we introduce a human feedback budget $H$, which sets a cap on the number of QA pairs that can be reviewed. When the number of identified high-conflict samples exceeds $H$, we select the top-$H$ examples with the highest conflict scores to prioritize feedback. This ensures annotation resources are directed toward the most necessary samples. Once human feedback is collected, we can fine-tune the proxy reward using any standard reward learning method (e.g., pairwise preference modeling). We then fine-tune the base policy using this improved reward model on any RM-based alignment algorithm (e.g., PPO). This loop can be repeated for multiple iterations, progressively refining both the reward and policy. We note that our conflict-aware sampling strategy is conceptually related to active learning and disagreement sampling (danka2018modal), but we believe it is underexplored in the context of proxy-policy misalignment in RLHF. A more thorough discussion of these connections is provided in the Appendix. Our algorithm also shares conceptual similarities with the Rejection Sampling (RSO) approach proposed by (liu2023statistical), as both involve sampling responses from the base policy and filtering them based on a proxy reward. However, our method differs from RSO, which is detailed in the Appendix. We also provide a detailed analysis of the scalability and computational efficiency of SHF-CAS, along with comparisons to PPO and RSO, in the Appendix.
@@
Tasks. To better understand the intrinsic conflicts between the base policy and the proxy reward, as discussed in the Method section, and to evaluate the effectiveness of our proposed SHF-CAS framework, we explore two fundamental alignment tasks: safety alignment and helpfulness alignment. For the safety alignment task, the goal is to fine-tune the LLM to avoid generating harmful content (dai2023safe; liu2024enhancing). We use the PKU-SafeRLHF dataset (dai2023safe), a high-quality human preference dataset designed to evaluate safety alignment in LLMs. Each example in the dataset includes a prompt, two candidate responses, a binary human preference label indicating which response is safer, and metadata identifying one or more of 19 harm categories. The dataset contains over 82.1K examples in total, with 73.9K for training and 8.2K held out for testing. For the helpfulness alignment task, the objective is to align LLMs to generate more helpful and informative responses to user queries (bai2022training; askell2021general). We conduct experiments using the Anthropic HH-RLHF helpful dataset (bai2022training). This dataset consists of prompts paired with two model-generated responses, along with human preference annotations indicating which response is more helpful. The prompts cover a wide range of general-purpose dialogue and instruction-following tasks. The dataset includes 161K comparisons for training and 8.5k for testing, collected from a human preference pipeline designed to train and evaluate helpfulness alignment in LLMs
@@
Setup. To emulate a realistic setting, we start with a strong pre-trained base policy $\pi_{\text{base}}$ and a biased proxy reward model $r_{\text{proxy}}$. For the safety alignmnet task, we construct the base policy by supervised fine-tuning a Pythia-6.9B model (biderman2023pythia) on the PKU-SafeRLHF training set, filtered to include 7 harm categories. This simulates a realistic scenario where the base policy is already strong on general safety objectives but lacks coverage for some specific harm types. To construct a less powerful and potentially biased proxy reward, we fine-tune a Pythia-1B model using the reward modeling procedure described in the Preliminaries Section. We restrict training to another 8 harm categories with 1 overlap category. This setup reflects a common real-world case where the proxy reward model is trained on limited supervision and biased toward certain categories. For the helpfulness alignment task, we follow a similar setup: the base policy is obtained by fine-tuning the Pythia-6.9B model on 30% of the Anthropic HH-RLHF training set, while the reward model is trained by fine-tuning the Pythia-1B model on a separate 30% split, with a 10% overlap in prompts between the two training sets to simulate knowledge known by both model.
@@
To identify proxy-policy conflicts, for both tasks, we sample responses from the entire training set. As part of our ablation study, we investigate the impact of key hyperparameters in our method, including the number of sampled responses per prompt $N$, the Kendall-Tau Distance threshold $\tau$, and the PACS thresholds $\delta$. To simulate oracle supervision in our experiments, we incorporate both model-based supervision and GPT-4o-based feedback. For model-based supervision, we adopt two high-quality reward models. For the safety alignment task, we use the beaver-7b-unified-cost model (dai2023safe), which is trained on the full PKU-SafeRLHF dataset and provides reward signals reflecting human safety preferences. For the helpfulness alignment task, we utilize RM-Mistral-7B (xiong2023iterative; dong2023raft), a reward model trained on a mixture of open-source helpfulness datasets, which has demonstrated strong alignment with human judgments in prior work. While human feedback remains the gold standard for alignment supervision, it is often impractical due to the high cost and latency of human annotation. Therefore, following the methodology of (chiang2023can), we additionally employ GPT-4o as an automated preference evaluator to approximate human supervision. This setup allows us to study alignment in more realistic settings where supervision may be noisy, biased, or resource-constrained, providing a more comprehensive evaluation of our method’s robustness. We then fine-tune the proxy reward on these high-conflict QA pairs using the same setup as the initial reward modeling process. For RM-based alignment, we fine-tune the base policy on the refined reward using PPO (schulman2017proximal). In these experiments, we assume one iteration of refinement and unlimited human feedback budget to clearly illustrate the impact of different parameters on alignment performance.
@@
As for baselines, we fine-tune the base policy using the vanilla PPO with the original proxy reward without incorporating any additional human feedback. In addition, we include the RSO (liu2023statistical), which similarly samples responses and filters them based on the proxy reward. For the RSO sampling, we adopt identical sampling parameters as our algorithm and use the hyperparemeters consistent with the settings in the original paper. For ablation analysis, we further include a variant in which the base policy is fine-tuned using PPO on a set of randomly sampled QA pairs, matched in number to those selected by our proposed method under each experimental configuration. This allows us to isolate the impact of conflict-aware sampling from the overall data quantity. Furthermore, to assess the benefits of multiple iterations, we iteratively apply the vanilla PPO, our SHF-CAS, and RSO algorithm for $i\in\{1,2,3\}$ iterations. More details about the experimental setup can be found in the Appendix.
@@
Evaluation. To ensure a fair comparison, we evaluate all models on the test set. We report the average reward obtained from the gold reward model to assess overall safety alignment. In addition, we compute the average PACS and Kendall-Tau Distance between each model and the gold reward to quantify the extent of residual proxy-policy misalignment after fine-tuning. For a more reliable and human-aligned evaluation, we further include GPT-4o (achiam2023gpt) as an automated judge. GPT-4o is used to compare outputs from two models for the same prompt and indicate a preference based on harmlessness or helpfulness. From these comparisons, we compute pairwise win-rates between all models. Additional details on the evaluation can be found in the Appendix.

### v2

@@ Introduction
Large Language Models (LLMs) have been undergoing rapid evolution (Guo et al. 2025; Pichai and Hassabis 2024; Achiam et al. 2023), and have demonstrated remarkable capabilities across a wide range of tasks, including following instructions (Chung et al. 2024; Ouyang et al. 2022; Liu et al. 2025), performing complex reasoning (Anil et al. 2023; Wei et al. 2022), and code generation (Gao et al. 2023; Ni et al. 2024). A key method for aligning these models with human intent is to fine-tune them using reward signals. A standard paradigm for this is Reinforcement Learning from Human Feedback (RLHF), where a pre-trained base policy $\pi_{\text{base}}$ is optimized using a reward model $r$ that approximates human preferences across dimensions such as helpfulness, safety, or formatting (Christiano et al. 2017; Chen et al. 2024; Ziegler et al. 2019). However, this approach rests on the assumption that the reward model accurately captures the true supervision signal. In practice, this assumption is frequently violated due to annotation noise, limited or biased human data (Jeon et al. 2020; Sadigh et al. 2017). As a result, the learned reward function becomes an imperfect proxy ($r_{\text{proxy}}$) of the true objective, which can lead to unintended behavior of the policy or sometimes serious consequences, such as reward over-optimization (Coste et al. 2023; Moskovitz et al. 2023) or reward hacking (Laidlaw et al. 2024; Wang et al. 2025; Denison et al. 2024; Everitt et al. 2021; Pan et al. 2022), where models optimize the proxy reward while diverging significantly from the actual intent.
In this work, we investigate how such misalignments manifest in practice and how we might systematically surface and mitigate their effects. Specifically, we consider a realistic scenario where we start from a strong base model $\pi_{\text{base}}$, already proficient on some tasks, and fine-tune it using a proxy reward $r_{\text{proxy}}$ trained on noisy or limited preference data from a related but distinct task. This setup is particularly relevant in practical settings where powerful models like GPT-4 (Achiam et al. 2023) already demonstrate strong general capabilities, and fine-tuning is applied to adapt them to niche or sensitive objectives. In such cases, we want to avoid degrading the base model’s performance or introducing unintended behaviors due to flawed or imprecise proxy supervision. Prior work has attempted to mitigate these effects either by regularizing the fine-tuning process, such as through KL-penalty (Liu and others 2020), or by enhancing the proxy reward model itself (Shen et al. 2024; Ramé et al. 2024). While effective, these approaches often lack insight into why such solutions work or where exactly the proxy model may be going wrong.
@@
Aligning LLMs with human intent often follows the reward-model-based alignment paradigm, also known for RLHF (Christiano et al. 2017; Stiennon et al. 2020), which typically consists of three major stages: (1) supervised fine-tuning, (2) reward modeling, and (3) reinforcement learning.
@@
Reward Modeling: To emulate human preferences, a reward model $r(x,y)$ is trained to score responses given a prompt $x$. A standard approach is to learn from pairwise human comparisons using the Bradley-Terry (BT) model (Bradley and Terry 1952), which defines the probability of preferring response $y_{1}$ over $y_{2}$ as: $p^{*}(y_{1}\succ y_{2}|x)=\frac{\exp(r(x,y_{1}))}{\exp(r(x,y_{1}))+\exp(r(x,y_{2}))}.$ Here, $y_{1}\succ y_{2}$ denotes $y_{1}$ is preferred and $y_{2}$ is dispreferred. Given a dataset $D=\{x^{i},y_{\omega}^{i}\succ y_{l}^{i}\}_{i=1}^{N}$ consisting of prompt-response pairs with human-labeled preferences, the reward function can be trained by minimizing the following logistic loss:
@@
This objective is typically optimized using proximal policy optimization (PPO) (Schulman et al. 2017).
@@
Proxy-Policy Alignment Conflict Score (PACS). A well-known observation in RLHF is that the optimal policy $\pi_{r}$ under a reward function $r$ for Equation 2 satisfies (Peng et al. 2019; Peters and Schaal 2007; Rafailov et al. 2023):
@@
Kendall-Tau Distance (K-T Distance). While PACS provides a localized conflict score, we also seek to assess the overall alignment between the base model and the proxy reward across a distribution of completions. To this end, we adopt the Kendall-Tau Distance (Abdi 2007), a standard measure of disagreement between two ranked lists.
@@
Building on the earlier insights that proxy-policy conflicts serve as valuable signals for additional supervision, and leveraging the defined conflict detection metrics, we now propose a practical algorithm SHF-CAS: Selective Human-in-the-Loop Feedback via Conflict-Aware Sampling. The complete procedure is described in Algorithm 1. It starts with identifying candidate QA pairs that exhibit high conflicts using the proposed PACS and K-T Distance metrics. These selected examples are then prioritized for human feedback. In particular, to make this approach resource-efficient, we introduce a human feedback budget $H$, which sets a cap on the number of QA pairs that can be reviewed. When the number of identified high-conflict samples exceeds $H$, we select the top-$H$ examples with the highest conflict scores to prioritize feedback. This ensures annotation resources are directed toward the most necessary samples. Once human feedback is collected, we can fine-tune the proxy reward using any standard reward learning method (e.g., pairwise preference modeling). We then fine-tune the base policy using this improved reward model on any RM-based alignment algorithm (e.g., PPO). This loop can be repeated for multiple iterations, progressively refining both the reward and policy.
@@
Tasks. To better understand the intrinsic conflicts between the base policy and the proxy reward, as discussed in the Method section, and to evaluate the effectiveness of our proposed SHF-CAS framework, we explore two fundamental alignment tasks: safety alignment and helpfulness alignment. For the safety alignment task, the goal is to fine-tune the LLM to avoid generating harmful content (Dai et al. 2023; Liu et al. 2024). We use the PKU-SafeRLHF dataset (Dai et al. 2023), a high-quality human preference dataset designed to evaluate safety alignment in LLMs. Each example in the dataset includes a prompt, two candidate responses, a binary human preference label indicating which response is safer, and metadata identifying one or more of 19 harm categories. The dataset contains over 82.1K examples in total, with 73.9K for training and 8.2K held out for testing. For the helpfulness alignment task, the objective is to align LLMs to generate more helpful and informative responses to user queries (Bai et al. 2022; Askell et al. 2021). We conduct experiments using the Anthropic HH-RLHF helpful dataset (Bai et al. 2022). This dataset consists of prompts paired with two model-generated responses, along with human preference annotations indicating which response is more helpful. The prompts cover a wide range of general-purpose dialogue and instruction-following tasks. The dataset includes 161K comparisons for training and 8.5k for testing, collected from a human preference pipeline designed to train and evaluate helpfulness alignment in LLMs
@@
Setup. To emulate a realistic setting, we start with a strong pre-trained base policy $\pi_{\text{base}}$ and a biased proxy reward model $r_{\text{proxy}}$. For the safety alignment task, we construct the base policy by supervised fine-tuning a Pythia-6.9B model (Biderman et al. 2023) on the PKU-SafeRLHF training set, filtered to include 7 harm categories. This simulates a realistic scenario where the base policy is already strong on general safety objectives but lacks coverage for some specific harm types. To construct a less powerful and potentially biased proxy reward, we fine-tune a Pythia-1B model using the reward modeling procedure described in the Preliminaries Section. We restrict training to another 8 harm categories with 1 overlap category. This setup reflects a common real-world case where the proxy reward model is trained on limited supervision and biased toward certain categories. For the helpfulness alignment task, we follow a similar setup: the base policy is obtained by fine-tuning the Pythia-6.9B model on 30% of the Anthropic HH-RLHF training set, while the reward model is trained by fine-tuning the Pythia-1B model on a separate 30% split, with a 10% overlap in prompts between the two training sets to simulate knowledge known by both model.
@@
To identify proxy-policy conflicts, for both tasks, we sample responses from the entire training set. As part of our ablation study, we investigate the impact of key hyperparameters in our method, including the number of sampled responses per prompt $N$, the Kendall-Tau Distance threshold $\tau$, and the PACS thresholds $\delta$. To simulate oracle supervision in our experiments, we incorporate both model-based supervision and GPT-4o-based feedback. For model-based supervision, we adopt two high-quality reward models. For the safety alignment task, we use the beaver-7b-unified-cost model (Dai et al. 2023), which is trained on the full PKU-SafeRLHF dataset and provides reward signals reflecting human safety preferences. For the helpfulness alignment task, we utilize RM-Mistral-7B (Xiong et al. 2023; Dong et al. 2023), a reward model trained on a mixture of open-source helpfulness datasets, which has demonstrated strong alignment with human judgments in prior work. While human feedback remains the gold standard for alignment supervision, it is often impractical due to the high cost and latency of human annotation. Therefore, following the methodology of (Chiang and Lee 2023), we additionally employ GPT-4o as an automated preference evaluator to approximate human supervision. This setup allows us to study alignment in more realistic settings where supervision may be noisy, biased, or resource-constrained, providing a more comprehensive evaluation of our method’s robustness. We then fine-tune the proxy reward on these high-conflict QA pairs using the same setup as the initial reward modeling process. For RM-based alignment, we fine-tune the base policy on the refined reward using PPO (Schulman et al. 2017). In these experiments, we assume one iteration of refinement and unlimited human feedback budget to clearly illustrate the impact of different parameters on alignment performance.
@@
As for baselines, we fine-tune the base policy using the vanilla PPO with the original proxy reward without incorporating any additional human feedback. In addition, we include the RSO (Liu et al. 2023), which similarly samples responses and filters them based on the proxy reward. For the RSO sampling, we adopt identical sampling parameters as our algorithm and use the hyperparameters consistent with the settings in the original paper. For ablation analysis, we further include a variant in which the base policy is fine-tuned using PPO on a set of randomly sampled QA pairs, matched in number to those selected by our proposed method under each experimental configuration. This allows us to isolate the impact of conflict-aware sampling from the overall data quantity. Furthermore, to assess the benefits of multiple iterations, we iteratively apply the vanilla PPO, our SHF-CAS, and RSO algorithm for $i\in\{1,2,3\}$ iterations. More details about the experimental setup can be found in the Appendix.
@@
Evaluation. To ensure a fair comparison, we evaluate all models on the test set. We report the average reward obtained from the gold reward model to assess overall safety alignment. In addition, we compute the average PACS and Kendall-Tau Distance between each model and the gold reward to quantify the extent of residual proxy-policy misalignment after fine-tuning. For a more reliable and human-aligned evaluation, we further include GPT-4o (Achiam et al. 2023) as an automated judge. GPT-4o is used to compare outputs from two models for the same prompt and indicate a preference based on harmlessness or helpfulness. From these comparisons, we compute pairwise win-rates between all models. Additional details on the evaluation can be found in the Appendix.
@@ Results
Conclusion
In this work, we investigate the challenges of aligning LLMs using reward-model-based fine-tuning under imperfect supervision. We propose that proxy-policy conflicts, disagreements between the base policy and the proxy reward, serve as useful indicators of potential misalignment, particularly in regions where both models may lack sufficient knowledge. To quantify such conflicts, we introduce two complementary metrics: the localized PACS and the global K-T Distance. Building on this insight, we design a practical algorithm named SHF-CAS, which prioritizes high-conflict QA pairs for external supervision. Experiments on two different alignment tasks highlight the potential of conflict-aware samplings to guide targeted feedback collection and improve alignment efficiency.